OpenStudy (anonymous):

Help! How would I solve this? √ 30*20*0.75

OpenStudy (anonymous):

That's suppose to be a long radical sign

OpenStudy (anonymous):

ok

OpenStudy (welshfella):

just grab your calculator and enter them in

OpenStudy (anonymous):

why don't you multiply the numbers inside first?

OpenStudy (anonymous):

lol ^^^

OpenStudy (welshfella):

thats a good way forward then take the square root of the result

OpenStudy (anonymous):

do you need a exact answer?

OpenStudy (anonymous):

450

OpenStudy (anonymous):

I'm trying to figure out how I would solve it

OpenStudy (anonymous):

450 ^(.5)

OpenStudy (anonymous):

or \[\sqrt{450}\]

OpenStudy (welshfella):

thats easy using a calculator

OpenStudy (anonymous):

It has to do with how fast a vehicle is going.

OpenStudy (welshfella):

or if you want an exact answer find the prime factors of 450

OpenStudy (anonymous):

Around 21?

OpenStudy (anonymous):

Is that how I would do it every single time just multiply and find the square root of that number?

OpenStudy (welshfella):

then pick out perfect squares from these factors

OpenStudy (welshfella):

yes - the calculator will give you a decimal answer that you acn correct to the number of decimals you may want

OpenStudy (welshfella):

21 is correct to the nearest whole number

OpenStudy (anonymous):

Ok great. Thank you so much for the help

OpenStudy (welshfella):

yw

OpenStudy (phi):

if you are doing it "by hand" \[ \sqrt{30 \cdot 20 \cdot 0.75 } \\ \frac{\sqrt{3\cdot 2\cdot 5 \cdot 2 \cdot 2 \cdot 5 \cdot 3 } }{\sqrt{4}} \]and look for pairs of numbers

OpenStudy (phi):

you would pull out a pair of 3's , a pair of 2's and a pair of 5's (leaving one 2 inside the square root \[ \frac{3\cdot 2 \cdot 5}{2} \sqrt{2} \] or \[ 15 \sqrt{2} \]